OpenStudy (anonymous):

Eliminate the parameter. x = radical t , y = 2t + 5

OpenStudy (anonymous):

I just started learning this, so you have\[x=\sqrt{t}\]and\[y=2t+5\]So you can get t by itself in x=sqrt t\[x=\sqrt{t}=>x ^{2}=t\]Plug t into y=2t+5\[y=2(x ^{2})+5\]

OpenStudy (anonymous):

so you just get t by itself and plug it in?

OpenStudy (anonymous):

I think so, and when graphed, you have to make sure only graph where t exists.

OpenStudy (anonymous):

Thank you! That was pretty easy haha also lets say i have an x=7t, for my next question, would the x be x=x/7?

OpenStudy (anonymous):

Did it just give you x=7t?

OpenStudy (anonymous):

No other equation?

OpenStudy (anonymous):

no x=7t, y=t+4 but i think the answer is y= x/7 +4 is that right?

OpenStudy (anonymous):

Yeah